### v1

# An iterative algorithm for the square-root Lasso

Abstract
In the framework of sparsity-enforcing regularisation for linear inverse problems, we consider the minimisation of a square-root Lasso cost function. We show that a simple modification of the Iterative Soft-Thresholding Algorithm (ISTA) for the Lasso problem leads to a minimiser of this cost function and we investigate the convergence properties of this algorithm. We also consider the group square-root Lasso, where sparsity is enforced for groups of variables instead of individual ones.
1 Introduction
We consider the following linear inverse problem: find an object of interest $f$ from the observation or measurement of $g=Af$ where $A$ is a given linear operator from a space $\mathcal{F}$, the object space, to which $f$ belongs, into a data or image space $\mathcal{G}$, to which $g$ belongs. The spaces $\mathcal{F}$ and $\mathcal{G}$ are assumed to be Hilbert spaces (typically $L^{2}$ spaces) and $A:\mathcal{F}\to\mathcal{G}$ to be a bounded linear operator. In the special instance of a finite-dimensional problem, $A$ is just a matrix and $\mathcal{F}$ and $\mathcal{G}$ Euclidean spaces. For simplicity, we will only consider real-valued $f$ and $g$.
Since measurements are affected by noise and the corresponding inverse problem is ill-posed, it should be regularised. A usual strategy to achieve this is to replace it by a variational problem. In the framework of sparsity-enforcing regularisation, the specific variational problem we consider in the present work is to find a minimiser $f_{*}$ for the following cost function
$\Phi_{\mu}(f)=\|Af-g\|+\mu\,\|f\|_{1}$ (1)
where $\mu>0$ is a fixed regularisation parameter and $\|f\|_{1}$ is a sparsity-enforcing norm defined by
$\|f\|_{1}=\sum_{\gamma\in\Gamma}w_{\gamma}\,|f_{\gamma}|.$ (2)
By $f_{\gamma}$ we denote the scalar product $f_{\gamma}=(f,\varphi_{\gamma})$ where $\{\varphi_{\gamma}\}_{\gamma\in\Gamma}$ is a fixed – but otherwise arbitrary – orthonormal basis of the object space $\mathcal{F}$. The “discrepancy” or “residual” $\|Af-g\|$ is measured by means of the norm of the image/data Hilbert space $\mathcal{G}$. It is well known that if $f$ is expected to be sparse in the given basis, i.e. if many coefficients $f_{\gamma}$ are expected to be zero, the penalty (2), which is a weighted $\ell^{1}$-norm on the sequence of coefficients $f_{\gamma}$ on that basis, will favour the recovery of such sparse objects. For simplicity, we will consider the case of unit weights $w_{\gamma}=1$ but this is not restrictive; the extension to arbitrary positive weights uniformly bounded from below $(w_{\gamma}\geq c>0)$ is straightforward and is needed when dealing with wavelet bases.
Dating back to [1], sparsity-enforcing regularisation has been abundantly studied in the literature but mostly resorting as in [1] to the minimisation of the following cost function
$\Phi_{\mu}^{\rm Lasso}(f)=\|Af-g\|^{2}+\tilde{\mu}\,\|f\|_{1}$ (3)
often referred to as the “Lasso” cost function. We notice that this latter cost function is not scale-invariant, in the sense that if $g$ is rescaled as $sg$ with a scale parameter $s>0$, as e.g. by a change of measurement unit, the corresponding minimiser of (3) is not $sf$. This may appear as a drawback of the Lasso cost function, which looks somewhat awkward from a dimensional point of view. The cost function (1), however, is scale invariant as is the usual Tikhonov cost function where the sparsity-enforcing penalty (2) in (3) is replaced by the quadratic penalty $\|f\|^{2}$. The minimisation of (1) is usually called the “square-root Lasso”. While its scale-invariance appears as an advantage, taking the square-root of the quadratic discrepancy renders it nonsmooth – as is the $\ell^{1}$-norm penalty (2) – and hence makes the optimisation problem more difficult.
We should remark though that both the Lasso and the square-root Lasso are convex problems. Hence they should be equivalent in the sense that a minimiser of (1) for a given value of $\mu$ should be a minimiser of (3) for some value of $\tilde{\mu}$. Also equivalent are the two associated constrained problems: minimise the discrepancy subject to an upper bound on the norm (2) or minimise the norm (2) subject to an upper bound on the discrepancy. However, the correspondence between the parameters involved (upper bounds and regularisation parameters) is not in general explicit and might be hard to compute. We will nevertheless show later in the paper that, for a given minimiser $f^{*}$ the correspondence between the respective square-root Lasso and Lasso parameters $\mu$ and $\tilde{\mu}$ is $\tilde{\mu}=2\mu\|Af^{*}-g\|$.
As the acronym Lasso (for Least absolute shrinkage and selection operator) created by Tibshirani [2], the appellation “square-root Lasso” (abbreviated as SQRT-Lasso or SR-Lasso) originated in statistics in a paper by Belloni et al. [3], in the framework of (finite-dimensional) linear regression, where (1) usually written in the traditional notations of the field as
$\Phi_{\mu}(\beta)=\|X\beta-y\|+\mu\,\|\beta\|_{1}.$ (4)
$X$ denotes the observed design matrix, $y$ the observed response and $\beta$ the unknown regression coefficient. The $\ell^{1}$-norm penalty on $\beta$, $\|\beta\|_{1}=\sum_{j}|\beta_{j}|$, enforces its sparsity, allowing for variable selection. In that seminal paper, the method is presented as a remedy to the fact that consistency results for the Lasso require a choice of the regularisation parameter $\mu$ that depends on an “oracle” knowledge of the variance of the noise (variance which is also hard to estimate). Indeed, the authors show that the square-root Lasso yields a “pivotal” estimator, in the sense that the same consistency results as for the Lasso can be obtained with a choice of $\mu$ which is independent of the variance of the noise.
After [3] the square-root Lasso has been extensively studied in the (high-dimensional) statistics literature [4, 5, 6, 7, 8, 9], and the results have also been extended to the group square-root Lasso [10], which similarly to the group Lasso [11, 8, 5] uses a penalty which favours the sparsity of groups of variables, i.e. the sum of the Euclidean norms (not squared) of each group. Its advantages have also been investigated in the framework of “Compressed Sensing” [12, 13]. We will not comment further on the pivotal properties derived in such settings since they require hypotheses, such as e.g. the restricted isometry property (RIP), the null-space property (NSP) or the restricted eigenvalue property (RE), which do not generally hold for the case of inverse imaging problems.
We believe that the simple algorithm we analyse here in the framework of inverse problems can be of interest not only for imaging problems (such as e.g. neuroimaging [14]) but also for these applications in statistics and compressed sensing. Because of the equivalence explained above with the Lasso, we do not of course expect an improvement in the quality of the reconstructions obtained by the square-root Lasso compared to the Lasso. However, the choice of the regularisation parameter could be eased and be less dependent on the effective noise level.
To the best of our knowledge, our algorithm differs from the other algorithmic solutions proposed in the literature: conic programming [3], alternating direction method of multipliers (ADMM) [15], coordinate descent [16, 17], proximal gradient descent and proximal Newton [18], and iteratively reweighted least squares (IRLS) [13]. An exception is the paper [10] where an algorithm similar to ours is proposed for the group square-root Lasso and convergence of the sequence of iterates is claimed. However, unfortunately, the proof of convergence provided in the Appendix of that paper appears to be incorrect for the reasons that will be explained in the following.
In Section 2, we introduce the proposed iterative algorithm, establishing a connection with the well-known “Iterative Soft-Thresholding Algorithm” (ISTA) for Lasso. Then we study its convergence properties and prove that the sequence of the iterates converges to a minimiser of the cost function (1) in finite-dimensional situations when the minimiser is unique. In Section 3, we generalise the algorithm to the case of group square-root Lasso and we introduce an iteratively reweighted least squares method to deal with the penalty. We also put a safeguard against the possibility of vanishing weights and we extend to this case the proof of convergence of Section 2.
2 A simple iterative algorithm for the square-root Lasso
2.1 Derivation of the algorithm
Notice that
${\rm argmin}_{f}\Phi_{\mu}(f)={\rm argmin}_{f,\sigma>0}\Phi_{\mu}(f,\sigma)$ (5)
where
$\Phi_{\mu}(f,\sigma)=\frac{\|Af-g\|^{2}}{\sigma}+\sigma+2\mu\,\|f\|_{1}.$ (6)
This trick is well known and has been used repeatedly in the literature. Notice that the cost function (6) is jointly convex in $f$ and $\sigma>0$ (see e.g. [19], Ex. 3.18).
This suggests the following alternating minimisation, starting from an arbitrary initial guess for $f_{0}$, and the corresponding residual $\sigma_{0}=\|Af_{0}-g\|$ (e.g. $f_{0}=0$ and $\sigma_{0}=\|g\|$), for $f$
$f_{k+1}={\rm argmin}_{f}\left[\frac{\|Af-g\|^{2}}{\sigma_{k}}+\sigma_{k}+2\mu\,\|f\|_{1}\right],$ (7)
or equivalently
$f_{k+1}={\rm argmin}_{f}\left[\|Af-g\|^{2}+2\mu\sigma_{k}\,\|f\|_{1}\right],\quad k=0,1,\dots,$ (8)
and for the residual
$\sigma_{k+1}={\rm argmin}_{\sigma>0}\left[\frac{\|Af_{k+1}-g\|^{2}}{\sigma}+\sigma+2\mu\,\|f_{k+1}\|_{1}\right],$ (9)
which simply yields
$\sigma_{k+1}=\|Af_{k+1}-g\|.$ (10)
Now the idea is to replace the minimisation (7), namely of $\Phi_{\mu}(f,\sigma_{k})$ with respect to $f$, which is a Lasso problem, by the minimisation of the surrogate cost function
${\Phi}^{\scriptstyle\rm{SUR}}_{\mu}(f,\sigma_{k};f_{k})=\Phi_{\mu}(f,\sigma_{k})+\frac{1}{\sigma_{k}}\left[\frac{1}{\tau}\|f-f_{k}\|^{2}-\|Af-Af_{k}\|^{2}\right]$ (11)
where $\tau$ is a positive parameter. If $\|A^{*}A\|=\|A\|^{2}\leq 1/\tau$ where $\|A\|$ denotes the operator spectral norm, this surrogate clearly satisfies the majorisation property
$\Phi_{\mu}(f,\sigma_{k})\leq{\Phi}^{\scriptstyle\rm{SUR}}_{\mu}(f,\sigma_{k};f_{k})~{},~{}~{}~{}~{}\mbox{{\rm for all}}~{}f,\sigma_{k}~{}\mbox{{\rm and}}~{}f_{k}.$ (12)
It also satisfies the anchoring property
$\Phi_{\mu}(f_{k},\sigma_{k})={\Phi}^{\scriptstyle\rm{SUR}}_{\mu}(f_{k},\sigma_{k};f_{k})~{},~{}~{}\mbox{{\rm for all}}~{}f_{k}~{}\mbox{{\rm and}}~{}\sigma_{k}.$ (13)
The surrogate (11) has been used in [1] to derive the ISTA algorithm for the Lasso problem. Optimisation strategies of this type are usually referred to as MM optimisation (for Majorisation-Minimisation) (see e.g. [20]). Thanks to the decoupling of the components of $f$, the minimiser of (11) is easily calculated by hand and is given by
$f_{k+1}={S}_{2\tau\mu\sigma_{k}}\left(f_{k}+\tau A^{*}(g-Af_{k})\right),$ (14)
where ${S}_{t}$ denotes the (nonlinear) “soft-thresholding” operator’, acting component-wise, meaning that its action on $h=\sum_{\gamma}h_{\gamma}\varphi_{\gamma}$ is defined by ${S}_{t}(h)=\sum_{\gamma}{S}_{t}(h_{\gamma})\varphi_{\gamma}$, and where $S_{t}(x)$ is the soft-thresholding function defined as
$S_{t}(x)=\left\{\begin{array}[]{ccl}x+t/2{}&~{}\mbox{if}{}&x\leq-t/2\\ 0{}&~{}\mbox{if}{}&|x|<t/2\\ x-t/2{}&~{}\mbox{if}{}&x\geq t/2.\end{array}\right.$ (15)
The update on $\sigma$ will be as above given by (10). The positive parameter $\tau$, also referred to as a relaxation parameter, characterises the stepsize taken in the direction of the negative gradient of $\|Af-g\|^{2}$. This iterative algorithm is a simple modification of ISTA, with a threshold varying at each iteration according to the current value $\sigma_{k}$ of the residual. Indeed, for the Lasso cost function (3), if the soft-thresholding operator ${S}_{2\tau\mu\sigma_{k}}$ in (14) is replaced by ${S}_{\tau\tilde{\mu}}$, we get the ISTA algorithm, which can be shown to converge to a minimiser of (3) [1].
Let us remark that a MM-optimisation strategy does not necessarily guarantee the convergence of the sequence of iterates to a minimum point of the cost function. This requires additional work, specific for each case.
Hence, to summarise, in the next subsection, we will investigate the convergence of the iterative algorithm with iterates $(f_{k},\sigma_{k}),k=0,1,\dots,$ and the update steps (14) and (10), namely
$f_{k+1}={S}_{2\tau\mu\sigma_{k}}\!\left(f_{k}+\tau A^{*}(g-Af_{k})\right),$ (16)
$\sigma_{k+1}=\|Af_{k+1}-g\|,$ (17)
initialised with arbitrary $f_{0}$ and $\sigma_{0}>0$.
Remark 1
The algorithm (16, 17) can also be interpreted as a variable step-length proximal-gradient algorithm (Theorem 3.4 in [21]) applied to the minimisation of (1). As the gradient of the first term is $A^{*}(Af-g)/\|Af-g\|$, the proximal gradient algorithm for (1) takes the form
$f_{k+1}=S_{2\mu\gamma_{k}}\left(f_{k}-\gamma_{k}\frac{A^{*}(Af_{k}-g)}{\|Af_{k}-g\|}\right).$ (18)
Setting $\gamma_{k}=\tau\|Af_{k}-g\|$ one formally obtains algorithm (16, 17). However, the usual convergence theorem (Theorem 3.4 in [21]) cannot be invoked because the first term in (1) is not everywhere differentiable (not when $Af=g$), its gradient is not Lipschitz continuous and the stepsizes $\gamma_{k}$ depend on the iterates themselves.
On the other hand, the proximal gradient algorithm of [22] is related to problem (8) in the sense that it updates the penalty parameter in every step (as in the cost function in expression (8)). However, due to the dependence of $\sigma_{k}$ on $f_{k}$, the algorithm of [22] cannot either be applied to the problem discussed here.
Therefore, due to these features, a careful and separate convergence analysis of the algorithm is necessary.
2.2 Convergence properties
A first nice property, typical of MM-optimisation methods, results from the derivation of the algorithm through surrogates. Indeed, this ensures a monotonic decrease of the cost function $\Phi_{\mu}(f,\sigma)$ at each iteration and, therefore, ensures its convergence.
Proposition 1
If $\tau\leq 1/\|A\|^{2}$, we have along the iteration (16, 17)
$\Phi_{\mu}(f_{k+1},\sigma_{k+1})\leq\Phi_{\mu}(f_{k},\sigma_{k});\qquad k=0,1,\dots.$ (19)
Hence the nonnegative and nonincreasing sequence $\{\Phi_{\mu}(f_{k},\sigma_{k})\}$ has a limit, say $\Phi_{\mu,*}$.
Proof: We have
$\Phi_{\mu}(f_{k+1},\sigma_{k+1})\leq\Phi_{\mu}(f_{k+1},\sigma_{k})$ (20)
since $\sigma_{k+1}$ minimises $\Phi_{\mu}(f_{k+1},\sigma)$. Moreover
$\Phi_{\mu}(f_{k+1},\sigma_{k})\leq{\Phi}^{\scriptstyle\rm{SUR}}_{\mu}(f_{k+1},\sigma_{k};f_{k})\leq{\Phi}^{\scriptstyle\rm{SUR}}_{\mu}(f_{k},\sigma_{k};f_{k})=\Phi_{\mu}(f_{k},\sigma_{k})$ (21)
by the fact that $f_{k+1}$ minimises (11) and by the majorisation and anchoring properties (12) and (13) of the surrogate.
We will now establish convergence properties of the iterates in a slightly generalised setting. As shown in [21], the ISTA algorithm as introduced in [1] can be reinterpreted as a proximal gradient algorithm (see also the books [23, 24]). We no longer require that $\|A^{*}A\|<1/\tau$, hence losing the majorisation property (12) but, as we will see, this will allow to prove convergence under a more flexible assumption on the link between the stepsize $\tau$ along the negative gradient and the norm of $A$.
Let us recall the definition of the “prox” $\bar{f}$ of $h$ with respect to a cost function $\Pi(f)$ (here the penalty)
$\bar{f}={\rm argmin}_{f}\left[\frac{1}{2}\ \|f-h\|^{2}+\Pi(f)\right].$ (22)
It is a well-known fact that for $\Pi(f)=\mu\sigma\|f\|_{1}$ we have $\bar{f}=S_{2\mu\sigma}(h)$.
To analyse the convergence of the iterative scheme (16, 17), we first establish two lemmas. The first lemma will enable us to control the decrease of the cost function for two successive iterates.
Lemma 1
The following inequality holds true for any pair of vectors $f_{k+1}$ and $f_{k}$
$\|Af_{k+1}-g\|^{2}-\|Af_{k}-g\|^{2}-2(A^{*}(Af_{k}-g),f_{k+1}-f_{k})\leq\|A\|^{2}\|f_{k+1}-f_{k}\|^{2}.$ (23)
Proof: The inequality follows by noting that, using elementary manipulations, the left-hand side of (23) is equal to $\|A(f_{k+1}-f_{k})\|^{2}$.
The second lemma in standard in convex analysis and subdifferential calculus (see [23], Section 16.4).
Lemma 2
Let $\bar{f}=S_{2\mu\sigma}(h)={\rm argmin}_{f}\left(\|f-h\|^{2}+2\mu\sigma\|f\|_{1}\right)$ for a given $h$. Then for any $f$ we have
$\mu\sigma\|\bar{f}\|_{1}\leq\mu\sigma\|f\|_{1}-(h-\bar{f},f-\bar{f}).$ (24)
Proof: The optimality condition is that 00 should belong to the subdifferential of the cost function at the point $\bar{f}$ or else that
$2(h-\bar{f})\in\partial(2\mu\sigma\|\bar{f}\|_{1}).$ (25)
By definition of the subdifferential, this implies the inequality (24).
Combining these two lemmas, we can establish the monotonic decrease of the cost function along the iteration, generalising Proposition 1.
Proposition 2
Provided that $0<\tau\leq 2/\|A\|^{2}$, the sequence of values of the cost function is monotonically decreasing along the iteration (16, 17):
$\Phi_{\mu}(f_{k+1},\sigma_{k+1})\leq\Phi_{\mu}(f_{k},\sigma_{k});\qquad k=0,1,\dots.$ (26)
Hence the nonnegative and nonincreasing sequence $\{\Phi_{\mu}(f_{k},\sigma_{k})\}$ has a limit, say $\Phi_{\mu,*}$.
Proof: We consider the inequality (24) of Lemma 2 for $f=f_{k}$, $h=f_{k}+\tau A^{*}(g-Af_{k})$ and, according to (14), $\bar{f}=f_{k+1}$, namely
$\tau\mu\sigma_{k}\|f_{k+1}\|_{1}\leq\tau\mu\sigma_{k}\|f_{k}\|_{1}-(f_{k}+\tau A^{*}(g-Af_{k})-f_{k+1},f_{k}-f_{k+1}).$ (27)
Adding to this latter inequality – multiplied by $2$ and divided by $\tau$ – the inequality (23) of Lemma 1, then dividing by $\sigma_{k}>0$ and adding $\sigma_{k}$ to both sides, we get, using (6),
$\Phi_{\mu}(f_{k+1},\sigma_{k})\leq\Phi_{\mu}(f_{k},\sigma_{k})+\frac{1}{\sigma_{k}}\left(\|A\|^{2}-\frac{2}{\tau}\right)\|f_{k+1}-f_{k}\|^{2}.$ (28)
Using the fact that $\sigma_{k+1}$ minimises $\Phi_{\mu}(f_{k+1},\sigma)$, we finally get
$0\leq\Phi_{\mu}(f_{k+1},\sigma_{k+1})\leq\Phi_{\mu}(f_{k},\sigma_{k})+\frac{1}{\sigma_{k}}\left(\|A\|^{2}-\frac{2}{\tau}\right)\|f_{k+1}-f_{k}\|^{2}.$ (29)
Hence, provided the last term is not positive, we can deduce that
$\Phi_{\mu}(f_{k+1},\sigma_{k+1})\leq\Phi_{\mu}(f_{k},\sigma_{k})\leq\dots\leq\Phi_{\mu}(f_{0},\sigma_{0}).$ (30)
This in turn implies that the sequences of the iterates and of the discrepancies are bounded.
Proposition 3
Provided that $0<\tau\leq 2/\|A\|^{2}$, the sequence of iterates $\{(f_{k},\sigma_{k})\}$ is bounded.
Proof: We have by (6) and (30)
$\Phi_{\mu}(f_{k},\sigma_{k})=2\sigma_{k}+2\mu\|f_{k}\|_{1}\leq\Phi_{\mu}(f_{0},\sigma_{0}).$ (31)
Since $\mu$ is fixed, we see that $\|f_{k}\|_{1}$ is uniformly bounded in $k$. By the inequality
$\|f\|^{2}=\sum_{\gamma}f_{\gamma}^{2}\leq{\rm max}_{\gamma}\{|f_{\gamma}|\}\sum_{\gamma}|f_{\gamma}|\leq(\sum_{\gamma}|f_{\gamma}|)^{2}=\|f\|_{1}^{2},$ (32)
we see that the norms $\|f_{k}\|$ are bounded as well.
Notice that this remains true in the case of a weighted $\ell^{1}$-norm, since with $c$ the strictly positive lower bound of the weights, we have $\|f\|^{2}=\sum_{\gamma}f_{\gamma}^{2}\leq{\rm max}_{\gamma}\left\{\frac{w_{\gamma}}{w_{\gamma}^{2}}\ |f_{\gamma}|\right\}\sum_{\gamma}w_{\gamma}|f_{\gamma}|\leq\frac{1}{c^{2}}(\sum_{\gamma}w_{\gamma}|f_{\gamma}|)^{2}=\frac{1}{c^{2}}\|f\|_{1}^{2}$.
Moreover by (31), we see that the residuals $\sigma_{k}$ are also uniformly bounded.
Let us denote by $\bar{\sigma}$ an upper bound for all $\sigma_{k}$’s. Using these results, we can now establish the so-called property of asymptotic regularity of the iterative scheme. This means that the norm of the difference between two successive iterates tends to zero with the number of iterations.
Proposition 4
(“asymptotic regularity”) Provided that $0<\tau<2/\|A\|^{2}$, for two successive iterates $f_{k}$ and $f_{k+1}$, we have that $\|f_{k+1}-f_{k}\|\to 0$ for $k\to\infty$. For the corresponding discrepancies, we also have $|\sigma_{k+1}-\sigma_{k}|\to 0$.
Proof: Since $\sigma_{k}\leq\bar{\sigma}$, we have by (29)
$\frac{1}{\bar{\sigma}}\left(\frac{2}{\tau}-\|A\|^{2}\right)\|f_{k+1}-f_{k}\|^{2}$$\leq$ $\frac{1}{\sigma_{k}}\left(\frac{2}{\tau}-\|A\|^{2}\right)\|f_{k+1}-f_{k}\|^{2}$ (33)
$\leq$ $\Phi_{\mu}(f_{k},\sigma_{k})-\Phi_{\mu}(f_{k+1},\sigma_{k+1}).$
Hence, by a telescopic sum argument,
$\frac{1}{\bar{\sigma}}\left(\frac{2}{\tau}-\|A\|^{2}\right)\sum_{k=0}^{K}\|f_{k+1}-f_{k}\|^{2}$$\leq$ $\sum_{k=0}^{K}\left[\Phi_{\mu}(f_{k},\sigma_{k})-\Phi_{\mu}(f_{k+1},\sigma_{k+1})\right]$ (34)
$=$ $\Phi_{\mu}(f_{0},\sigma_{0})-\Phi_{\mu}(f_{K+1},\sigma_{K+1})$
$\leq$ $\Phi_{\mu}(f_{0},\sigma_{0}).$
From this we conclude that
$\sum_{k=0}^{\infty}\|f_{k+1}-f_{k}\|^{2}<+\infty,\ \ \hbox{implying}\ \ \|f_{k+1}-f_{k}\|\to 0\ \ \hbox{for}\ \ k\to\infty.$ (35)
Moreover,
$|\sigma_{k+1}-\sigma_{k}|$ $=$ $|\ \|Af_{k+1}-g\|-\|Af_{k}-g\|\ |$ (36)
$\leq$ $\|A(f_{k+1}-f_{k})\|\leq\|A\|\ \|f_{k+1}-f_{k}\|$
using the reverse triangular inequality. Hence the asymptotic regularity for the residual follows from that of the iterates $f_{k}$.
To proceed further, we exploit the fact that the sequence of iterates is bounded. Therefore, in finite dimension, it contains convergent subsequences. As we will see, this last property is crucial for our convergence proof. Accordingly, although the derivation up to this point would hold also in an infinite-dimensional Hilbert space setting, we will assume from now on that $f$ belongs to $\mathbb{R}^{d}$. We will also assume that $0<\tau<2/\|A\|^{2}$.
Let us denote by $T$ the combined iteration mapping, which maps the pair of iterates $(f_{k},\sigma_{k})$ onto the next one $(f_{k+1},\sigma_{k+1})$. According to (14) and (10), it is given by
$T(f,\sigma)=(\tilde{f},\tilde{\sigma})\ \hbox{with}$ $\tilde{f}$ $={S}_{2\tau\mu\sigma}\!\left(f+\tau A^{*}(g-Af)\right),$
[…]
We observe that $T$ is a continuous mapping in both $f$ and $\sigma$. Notice that contrary to the mapping applying $f_{k}$ on $f_{k+1}$, which depends on $k$ via the threshold $\sigma_{k}$, this combined mapping is stationary, i.e. does not depend on $k$. The next proposition shows that the limit $(f_{*},\sigma_{*})$ of any convergent subsequence of the (bounded) sequence of iterates is a fixed point of the mapping $T$, namely that $(f_{*},\sigma_{*})=T(f_{*},\sigma_{*})$.
Proposition 5
Let $\{(f_{k_{j}},\sigma_{k_{j}})\}$ be a convergent subsequence of the sequence of iterates $\{(f_{k},\sigma_{k})\}$, with limit $(f_{*},\sigma_{*})$. Then it converges to a fixed point of the iteration mapping $T$, i.e. $(f_{*},\sigma_{*})=T(f_{*},\sigma_{*})$, or equivalently
$f_{*}={S}_{2\tau\mu\sigma_{*}}\!\left(f_{*}+\tau A^{*}(g-Af_{*})\right)\ \hbox{with}\ \sigma_{*}=\|Af_{*}-g\|.$ (38)
Proof: If $(f_{k_{j}},\sigma_{k_{j}})\stackrel{{\scriptstyle j\to\infty}}{{\longrightarrow}}(f_{*},\sigma_{*})$ then, by Proposition 4, also $(f_{k_{j}+1},\sigma_{k_{j}+1})\stackrel{{\scriptstyle j\to\infty}}{{\longrightarrow}}(f_{*},\sigma_{*})$. By taking the same limit on both sides of the update rule $(f_{k_{j}+1},\sigma_{k_{j}+1})=T(f_{k_{j}},\sigma_{k_{j}})$, one finds, using the continuity of $T$, that $(f_{*},\sigma_{*})=T(f_{*},\sigma_{*})$.
The next step is to show that a fixed point of the mapping $T$ is a minimum point of the square-root Lasso cost function (1). Indeed, the following proposition holds true.
Proposition 6
A fixed point of the iteration $(f_{*},\sigma_{*})=T(f_{*},\sigma_{*})$ with $T$ given by (2.2) is a minimiser of the cost function (1).
Proof: The optimality conditions for (1), resulting from subdifferential calculus, are
$|A^{*}(g-Af_{*})|_{\gamma}\leq\mu\|g-Af_{*}\|\ \hbox{with equality if and onlyif}\ {(f_{*})}_{\gamma}\neq 0$ (39)
(i.e. if and only if $(f_{*})_{\gamma}$ is a so-called “active” component). These conditions are precisely equivalent to the fixed-point conditions.
Remark 2
Using the optimality conditions, we can see that, as already announced in the introduction, the minimisers of the Lasso cost function $\|Af-g\|^{2}+\tilde{\mu}\|f\|_{1}$ and the minimisers of the square-root Lasso cost function $\|Af-g\|+\mu\|f\|_{1}$ coincide when $\tilde{\mu}=2\mu\|Af_{*}-g\|$ since they respectively satisfy
$0\in 2A^{*}(Af_{*}-g)+\tilde{\mu}\partial\|f_{*}\|_{1}\ \hbox{and}\ 0\in A^{*}(Af_{*}-g)/\|Af_{*}-g\|+\mu\partial\|f_{*}\|_{1}$ (40)
when $Af_{*}\neq g$. Notice, however, that if the square-root Lasso has multiple solutions for a given $\mu$, then these may correspond to Lasso solutions for different values of $\tilde{\mu}$.
The previous results allow us in finite dimension to prove the convergence the sequence of iterates to a minimiser of the cost function provided that such minimiser is unique.
Theorem 1
If the minimiser of the square-root Lasso cost function (1) is unique, then the sequence of iterates $(f_{k},\sigma_{k})$ defined by (16, 17) converges to this minimiser.
Proof: By Proposition 3, we know that the iterates are bounded and by Propositions 5 and 6 that any limit point is a fixed point, which is also a minimum point of (1). If the minimiser is unique, the bounded sequence can have only a single limit point, and hence must converge to this minimiser.
Remark 3
Although it may appear as a strong restriction, let us remark that most often uniqueness will hold true. Indeed, to have more than one solution, the principal directions of $A$ would have to “conspire” with the position of the (weighted) $\ell^{1}$-ball in ${\mathbb{R}}^{d}$ in order to align with an edge or facet of such ball. Indeed, a minimiser of (1) has to be a tangency point between a level set of the discrepancy and of the (weighted) $\ell^{1}$-ball. Thus in general, nonuniqueness is unlikely to happen.
Remark 4
We can recall the following result about the limit points of an iteration with asymptotic regularity, known as Ostrowski’s theorem (see the book by Ostrowski [25] or Proposition 8.2.1 in the book by Lange [26]). However, this does not allow to establish the convergence to a minimum point.
Proposition 7
Let $f_{k}$ be a bounded sequence in $\mathbb{R}^{d}$. If, for some norm $\|f\|$,
$\lim_{k\rightarrow+\infty}\|f_{k+1}-f_{k}\|=0,$ (41)
then the set of the limit points of $f_{k}$ is a compact connected set. Furthermore, if this set is finite, then it reduces to a single point and the sequence converges.
3 Extension to the group square-root Lasso
The group Lasso aims at enforcing sparsity of groups of variables through a penalty that is the sum of the (quadratic) norms of each group. We assume that there is no overlap between the different groups of variables in $f$, which we denote by $f^{(j)},j=1,\dots,J$. Hence the cost function to minimise for the group square-root Lasso is given by
$\Phi^{\rm G}_{\mu}(f)=\|Af-g\|+\mu\,\sum_{j}\|f^{(j)}\|.$ (42)
For simplicity, we have used the same regularisation parameter $\mu$ for each group. This is not restrictive, however, and we could easily use a different parameter $\mu_{j}$ for each of the groups.
As above, we see that
${\rm argmin}_{f}\Phi^{\rm G}_{\mu}(f)={\rm argmin}_{f,\sigma>0}\Phi^{\rm G}_{\mu}(f,\sigma)$ (43)
where
$\Phi^{\rm G}_{\mu}(f,\sigma)=\frac{\|Af-g\|^{2}}{\sigma}+\sigma+2\mu\,\sum_{j}\|f^{(j)}\|.$ (44)
Starting from arbitrary $\sigma_{0}$ and $f_{0}^{(j)}$’s, we can define the alternating iteration for $f$
$f_{k+1}={\rm argmin}_{f}\left[\|Af-g\|^{2}+2\mu\sigma_{k}\,\sum_{j}\|f^{(j)}\|\right],\quad k=0,1,\dots,$ (45)
and for the residual
$\sigma_{k+1}=\|Af_{k+1}-g\|,$ (46)
and form the surrogate cost function
${(\Phi^{\rm G}_{\mu})}^{\scriptstyle\rm{SUR}}(f,\sigma_{k};f_{k})=\Phi^{\rm G}_{\mu}(f,\sigma_{k})+\frac{1}{\sigma_{k}}\left[\frac{1}{\tau}\|f-f_{k}\|^{2}-\|Af-Af_{k}\|^{2}\right].$ (47)
In analogy with (14), its minimiser $f_{k+1}$ is given by replacing the soft-thresholding operator by a block soft-thresholding operator whose expression can be easily derived as a “prox” (see e.g. [21] or [24])
$S^{\rm G}_{t}(x)=\frac{x}{\|x\|}{\rm max}\left\{0,\|x\|-\frac{t}{2}\right\}$ (48)
to be applied on each group. This yields the minimiser
$f^{(j)}_{k+1}=S^{\rm G}_{2\tau\mu\sigma_{k}}\!\left(f^{(j)}_{k}+\tau[A^{*}(g-Af_{k})]^{(j)}\right)$ (49)
where $f^{(j)}$ denotes the vector formed by the components of the vector $f$ belonging to the group $j$.
The proofs of the previous section can be extended to this case in a straightforward manner, establishing convergence of the sequence of iterates to a minimiser of the cost function (42) in a finite-dimensional situation and provided that the minimiser is unique.
Remark 5
We must remark that the algorithm (46, 49) has been previously written in [10]. However, the proof of convergence provided there is not correct. Indeed, it critically relies on a fixed-point theorem for non-expansive mappings due to Opial [27]. This theorem had been used in [1] to prove convergence of the ISTA algorithm for Lasso (see also there for a simplification of Opial’s proof). In the case of ISTA, the iteration map $T$ is indeed non-expansive since it is the product of the non-expansive Landweber operator $Lf=f+A^{*}(g-Af)$ with the soft-thresholding operator ${S}_{t}$ with a fixed threshold $t$. However, in the case of (46, 49), the threshold varies along the iteration and depends on the residual $\sigma_{k}=\|Af_{k}-g\|$. Hence, with $\sigma(u)=\|Au-g\|$ and $\sigma(v)=\|Av-g\|$, $\|S^{\rm G}_{2\tau\mu\sigma(u)}(u)-S^{\rm G}_{2\tau\mu\sigma(v)}(v)\|$ cannot be bounded by $\|u-v\|$ as it is easy to see on a one-dimensional example. This unfortunately invalidates the proof of convergence given in the Appendix of [10].
An alternative to the previous iteration is obtained when using the same trick for the penalty as for the discrepancy and noticing that
${\rm argmin}_{f}\Phi^{\rm G}_{\mu}(f)={\rm argmin}_{f,\sigma>0,\rho^{(j)}>0}\Phi^{\rm G}_{\mu}(f,\sigma,\rho^{(1)},\dots,\rho^{(J)})$ (50)
with
$\Phi^{\rm G}_{\mu}(f,\sigma,\rho^{(1)},\dots,\rho^{(J)})=\frac{\|Af-g\|^{2}}{\sigma}+\sigma+\mu\,\sum_{j}\left(\frac{\|f^{(j)}\|^{2}}{\rho^{(j)}}+\rho^{(j)}\right).$ (51)
This cost function is jointly convex in $f$, $\sigma$ and the $\rho^{(j)}$’s. Starting from an arbitrary $f_{0}$ with subgroups of components $f_{0}^{(j)}\neq 0$ as well as from $\sigma_{0}=\|Af_{0}-g\|$ and $\rho^{(j)}_{0}=\|f_{0}^{(j)}\|$, we could in principle proceed with the following iteration
$f_{k+1}={\rm argmin}_{f}\left[\|Af-g\|^{2}+\mu\sigma_{k}\sum_{j}\frac{1}{\rho_{k}^{(j)}}\|f^{(j)}\|^{2}\right],\quad k=0,1,\dots,$ (52)
with the updates for the $\rho^{(j)}$’s
$\rho_{k+1}^{(j)}=\|f_{k+1}^{(j)}\|$ (53)
and for the residual
$\sigma_{k+1}=\|Af_{k+1}-g\|.$ (54)
The minimisation (52) then reduces to a quadratic Tikhonov’s method for each group of variables.
We could also avoid the corresponding matrix inversions by using again a surrogate analogous to (47). Then (49) will be replaced by the damped Landweber iteration
$f^{(j)}_{k+1}=\frac{1}{1+\tau\mu\frac{\sigma_{k}}{\rho_{k}^{(j)}}}\!\left(f^{(j)}_{k}+\tau[A^{*}(g-Af_{k})]^{(j)}\right),\ \ j=1,\dots,J,$ (55)
with a damping factor which is different for each group and varies along the iteration.
As done in [28] for the penalty term of the Lasso cost function, we have an instance of IRLS (Iteratively Reweighted Least Squares). Notice that in [13] IRLS is applied for both the penalty and the discrepancy of the square-root Lasso cost function.
However, the iteration (55) does not take into account the possibility of vanishing values of $\rho_{k}^{(j)}$, which is to be expected when enforcing group sparsity. We remark that, on the contrary, the possibility of a vanishing discrepancy $\sigma_{k}$ for some $k$ is quite unlikely. Would this nevertheless occur, the problem could be easily fixed by resetting $\sigma_{k}$ to some positive small value and consider this as a restart of the iteration. We need however a safeguard when dealing with this problem for the penalty. To this aim, we use a trick similar to the one used in [28]. Accordingly, let us modify (51) as follows and define
$\Phi^{\rm G}_{\mu}(f,\sigma,\varepsilon,\rho^{(1)},\dots,\rho^{(J)})$ $=$ $\frac{\|Af-g\|^{2}}{\sigma}+\sigma$ (56)
$+$ $\mu\,\sum_{j}\left(\frac{\|f^{(j)}\|^{2}+\varepsilon^{2}}{\rho^{(j)}}+\rho^{(j)}\right).$
To ensure convergence to a minimiser of the original cost function, the parameter $\varepsilon$ should tend to zero for $k\to\infty$. Therefore we introduce the two new update rules
$\varepsilon_{k+1}={\rm min}\{\varepsilon_{k},\sqrt{\|f_{k+1}-f_{k}\|}\}$ (57)
and
$\rho_{k+1}^{(j)}={\rm argmin}_{\rho^{(j)}}\left(\frac{\|f_{k+1}^{(j)}\|^{2}+\varepsilon_{k+1}^{2}}{\rho^{(j)}}+\rho^{(j)}\right)=\sqrt{\|f_{k+1}^{(j)}\|^{2}+\varepsilon_{k+1}^{2}},$ (58)
for $j=1,\dots,J$.
As for the square-root Lasso, we can now prove the monotonic decay of the cost function (56) along the iteration.
Proposition 8
Provided that $0<\tau\leq 2/\|A\|^{2}$, the sequence of values of the cost function (56) is monotonically decreasing along the iteration (55, 57, 58) , i.e. for $k=0,1,\dots$,
$\Phi^{\rm G}_{\mu}(f_{k+1},\sigma_{k+1},\varepsilon_{k+1},\rho_{k+1}^{(1)},\dots,\rho_{k+1}^{(J)})\leq\Phi^{\rm G}_{\mu}(f_{k},\sigma_{k},\varepsilon_{k},\rho_{k}^{(1)},\dots,\rho_{k}^{(J)}).$ (59)
Hence the nonnegative and nonincreasing sequence $\{(\Phi^{G}_{\mu}(f_{k},\sigma_{k},\varepsilon_{k},\rho_{k}^{(1)},\dots,\rho_{k}^{(J)})\}$ has a limit, say $\Phi^{G}_{\mu,*}$.
Proof: We consider the inequality (24) of Lemma 2 for $f=f_{k}$, $h=f_{k}+\tau A^{*}(g-Af_{k})$ and, in agreement with (55), $\bar{f}=f_{k+1}$, so that we get the following inequality, analogous to (27) for the new penalty,
$\tau\mu\sigma_{k}\sum_{j}\frac{\|f_{k+1}^{(j)}\|^{2}+\varepsilon_{k}^{2}}{\rho_{k}^{(j)}}$$\leq$ $\tau\mu\sigma_{k}\sum_{j}\frac{\|f_{k}^{(j)}\|^{2}+\varepsilon_{k}^{2}}{\rho_{k}^{(j)}}$ (60)
-- $2(f_{k}+\tau A^{*}(g-Af_{k})-f_{k+1},f_{k}-f_{k+1}).$
Adding to this inequality divided by $\tau$ the inequality (23) of Lemma 1, and then dividing by $\sigma_{k}>0$ and adding $\sigma_{k}+\mu\sum_{j}\rho_{k}^{(j)}$ to both sides, we get
$\Phi^{G}_{\mu}(f_{k+1},\sigma_{k},\varepsilon_{k},\rho_{k}^{(1)},\dots,\rho_{k}^{(J)})$$\leq$ $\Phi^{\rm G}_{\mu}(f_{k},\sigma_{k},\varepsilon_{k},\rho_{k}^{(1)},\dots,\rho_{k}^{(J)})$ (61)
$+$ $\frac{1}{\sigma_{k}}\left(\|A\|^{2}-\frac{2}{\tau}\right)\|f_{k+1}-f_{k}\|^{2}.$
Since, by (57), $\varepsilon_{k+1}\leq\varepsilon_{k}$, we have $\Phi^{G}_{\mu}(f_{k+1},\sigma_{k},\varepsilon_{k+1},\rho_{k}^{(1)},\dots,\rho_{k}^{(J)})\leq\Phi^{\rm G}_{\mu}(f_{k+1},\sigma_{k},\varepsilon_{k},\rho_{k}^{(1)},\dots,\rho_{k}^{(J)})$. Then using the update property (58) for the $\rho^{(j)}$’s and the fact that $\sigma_{k+1}$ minimises $\frac{\|Af_{k+1}-g\|^{2}}{\sigma}+\sigma$, we finally get
$\Phi^{\rm G}_{\mu}(f_{k+1},\sigma_{k+1},\varepsilon_{k+1},\rho_{k+1}^{(1)},\dots,\rho_{k+1}^{(J)})$$\leq$ $\Phi^{\rm G}_{\mu}(f_{k},\sigma_{k},\varepsilon_{k},\rho_{k}^{(1)},\dots,\rho_{k}^{(J)})$ (62)
$+$ $\frac{1}{\sigma_{k}}\left(\|A\|^{2}-\frac{2}{\tau}\right)\|f_{k+1}-f_{k}\|^{2}.$
Hence, provided the last term in this inequality is not positive, we can deduce that
$\Phi^{\rm G}_{\mu}(f_{k+1},\sigma_{k+1},\varepsilon_{k+1},\rho_{k+1}^{(1)},\dots,\rho_{k+1}^{(J)})$$\leq$ $\Phi^{\rm G}_{\mu}(f_{k},\sigma_{k},\varepsilon_{k},\rho_{k}^{(1)},\dots,\rho_{k}^{(J)})$ (63)
$\leq$ $\Phi^{\rm G}_{\mu}(f_{0},\sigma_{0},\varepsilon_{0},\rho_{0}^{(1)},\dots,\rho_{0}^{(J)}).$
Now, thanks to the latter proposition, we can establish the analogues of Propositions 3 and 4, namely the uniform boundedness of the iterates and their asymptotic regularity.
Proposition 9
Provided that $0<\tau\leq 2/\|A\|^{2}$, the sequence of iterates $\{(f_{k},\sigma_{k},\varepsilon_{k},\rho_{k}^{(1)},\dots,\rho_{k}^{(J)})\}$ is bounded.
Proof: Indeed we have
$\Phi^{\rm G}_{\mu}(f_{k},\sigma_{k},\varepsilon_{k},\rho_{k}^{(1)},\dots,\rho_{k}^{(J)})$ $=$ $2\sigma_{k}+2\mu\sum_{j}\rho_{k}^{(j)}$ (64)
$\leq$ $\Phi^{\rm G}_{\mu}(f_{0},\sigma_{0},\varepsilon_{0},\rho_{0}^{(1)},\dots,\rho_{0}^{(J)})$
so that, since $\mu$ is fixed, we see that the residuals $\sigma_{k}$ are uniformly bounded, say by $\bar{\sigma}$. Also $\rho_{k}^{(j)}=\sqrt{\|f^{(j)}_{k}\|^{2}+\varepsilon_{k}^{2}}$ is uniformly bounded in $k$ for any $j$. This implies that the norms $\|f^{(j)}_{k}\|$ and hence also the norms $\|f_{k}\|$ are uniformly bounded as well.
Proposition 10
(“asymptotic regularity”) Provided that $0<\tau<2/\|A\|^{2}$, for two successive iterates $f_{k}$ and $f_{k+1}$, we have that $\|f_{k+1}-f_{k}\|\to 0$ for $k\to\infty$. We also have $|\sigma_{k+1}-\sigma_{k}|\to 0$ and $|\rho_{k+1}^{(j)}-\rho_{k}^{(j)}|\to 0$.
Proof: Since $\sigma_{k}\leq\bar{\sigma}$, we have by (62)
$\frac{1}{\bar{\sigma}}\left(\frac{2}{\tau}-\|A\|^{2}\right)\|f_{k+1}-f_{k}\|^{2}\leq\frac{1}{\sigma_{k}}\left(\frac{2}{\tau}-\|A\|^{2}\right)\|f_{k+1}-f_{k}\|^{2}$
$\leq\Phi^{\rm G}_{\mu}(f_{k},\sigma_{k},\varepsilon_{k},\rho_{k}^{(1)},\dots,\rho_{k}^{(J)})-\Phi^{\rm G}_{\mu}(f_{k+1},\sigma_{k+1},\varepsilon_{k+1},\rho_{k+1}^{(1)},\dots,\rho_{k+1}^{(J)}).$ (65)
Hence, by Proposition 8 and the telescopic argument used in Proposition 4, we get
$\frac{1}{\bar{\sigma}}\left(\frac{2}{\tau}-\|A\|^{2}\right)\sum_{k=0}^{K}\|f_{k+1}-f_{k}\|^{2}\leq\Phi_{\mu}(f_{0},\sigma_{0},\varepsilon_{0},\rho_{0}^{(1)},\dots,\rho_{0}^{(J)}).$ (66)
From this we conclude that
$\sum_{k=0}^{\infty}\|f_{k+1}-f_{k}\|^{2}<+\infty,\ \ \hbox{implying}\ \ \|f_{k+1}-f_{k}\|\to 0\ \ \hbox{for}\ \ k\to\infty.$ (67)
Moreover, $|\varepsilon_{k+1}-\varepsilon_{k}|\to 0$ since $\varepsilon_{k}\to 0$, $|\sigma_{k+1}-\sigma_{k}|\to 0$ as in (36) and $|\rho_{k+1}^{(j)}-\rho_{k}^{(j)}|\to 0$ since $\rho_{k}^{(j)}=\|f_{k}^{(j)}\|$ in the limit $\varepsilon_{k}\to 0$ and
$|\;\|f_{k+1}^{(j)}\|-\|f_{k}^{(j)}\|\;|\leq\|f_{k+1}^{(j)}-f_{k}^{(j)}\|\leq\|f_{k+1}-f_{k}\|\to 0.$ (68)
by the reverse triangular inequality.
As in Section 2.2, we can now define the combined iteration mapping, which maps a set of iterates $(f_{k},\sigma_{k},\varepsilon_{k},\rho_{k}^{(1)},\dots,\rho_{k}^{(J)})$ onto the next one $(f_{k+1},\sigma_{k+1},\varepsilon_{k+1},\rho_{k+1}^{(1)},\dots,\rho_{k+1}^{(J)})$. According to (55) and the update rules (54), (57) and (58), it is given by
$T^{\rm G}(f,\sigma,\varepsilon,\rho^{(1)},\dots,\rho^{(J)})=(\tilde{f},\tilde{\sigma},\tilde{\varepsilon},\tilde{\rho}^{(1)},\dots,\tilde{\rho}^{(J)})\quad\hbox{with for}\ j=1\dots,J,$
${\tilde{f}}^{(j)}=\frac{1}{1+\tau\mu\frac{\sigma}{\rho^{(j)}}}\!\left(f^{(j)}+\tau[A^{*}(g-Af)]^{(j)}\right),$
$\tilde{\sigma}=\|A\tilde{f}-g\|,$
$\tilde{\varepsilon}={\rm min}\{\varepsilon,\sqrt{\|\tilde{f}-f\|}\},$
${\tilde{\rho}}^{(j)}=\sqrt{\|{\tilde{f}}^{(j)}\|^{2}+\tilde{\varepsilon}^{2}}.$ (69)
In finite dimension, with a similar reasoning as in Proposition 5, we can then show that the limit of any converging subsequence of the iterates is a fixed point of the iteration. Moreover, as in Proposition 6, checking that any fixed point of $T^{\rm G}$ satisfies the optimality conditions, we have that it converges to a minimum point. The proof is straightforward so we will not report it in detail. Summarising, if the cost function admits a unique minimiser, we have the following result.
Theorem 2
If the minimiser of the group square-root Lasso cost function (42) is unique, then the sequence of iterates converges to this unique minimiser.
Proof: By Proposition 9, we know that the iterates are bounded and by the analogues of Propositions 5 and 6 that any limit point is a fixed point, which is also a minimum point of (42). If the minimiser is unique, the bounded sequence can have only a single limit point, and hence must converge to this minimiser.
Dedication
The present paper is dedicated to the memory of Pierre Célestin Sabatier who played a major role in establishing the field of inverse problems as a recognised independent scientific domain. Besides, as the seed motivating this work, we acknowledge a discussion on the lack of scale invariance of the Lasso with Mario Bertero, to whom we also want to bear homage.